Tina was three times younger than Helen in 2006. In 2012 Helen will be 42. How old will Tina be in 2012?
Found 2 solutions by Alan3354, ankor@dixie-net.com:
Answer by Alan3354(69443)

Tina was three times younger than Helen in 2006. In 2012 Helen will be 42. How old will Tina be in 2012?
---------------
"three times younger" is nonsense.
1 time younger is zero.
-----
10 times closer, 100 times smaller, etc. are also gibberish

Answer by ankor@dixie-net.com(22740)

Tina was three times younger than Helen in 2006.
3t = h
In 2012 Helen will be 42.
h + 6 = 42
How old will Tina be in 2012?
:
Six years later
3t+6 = h+6
h+6=42 therefore
3t + 6 = 42
3t = 42 - 6
3t = 36
t = 36/12
t = 12 in 2006
Therefore 6 years in 2012, Tina will be 18